# Enhancing the Transferability of Adversarial Attacks on Face Recognition with Diverse Parameters Augmentation

### v1

@@ 6.5 Visual Quality Study
Furthermore, we evaluate the visual quality of our proposed method against that of previous attack methods. We choose FIM, DI, DFANet, VMI, SSA, SIA, BPFA, and BSR as comparative baselines and generate adversarial examples using MobileFace as the surrogate model on the LFW dataset. The experimental configuration is consistent with the one detailed in Tab. 3. The outcomes are depicted in Fig. 8. As shown in Fig. 8, our proposed method achieves visual quality performance on par with other methods. Notably, the transferability of the adversarial examples generated by our method significantly exceeds that of the baselines, which further underscores the superiority of our proposed attack method.

### v3

@@ 7.6 Visual Quality Study
Furthermore, we evaluate the visual quality of our proposed method against that of previous attack methods. We choose FIM, DI, DFANet, VMI, SSA, SIA, BPFA, and BSR as comparative baselines and generate adversarial examples using MobileFace as the surrogate model on the LFW dataset. The experimental configuration is consistent with the one detailed in Table 1. The outcomes are depicted in Figure 8. As shown in Figure 8, our proposed method achieves visual quality performance on par with other methods. Notably, the transferability of the adversarial examples generated by our method significantly exceeds that of the baselines, which further underscores the superiority of our proposed method.
7.7 Ethics and Potential Broader Impact
This paper introduces research that contributes to the advancement of the field within Computer Vision and Pattern Recognition. The attack method we propose poses a potential threat to the security of FR models. Our goal is to heighten awareness through this proposed method and strengthen the resilience of FR models against such vulnerabilities. We prohibit the use of the attack method presented in this paper for commercial purposes. Furthermore, we explicitly forbid the application of the proposed attack method in any attempt to compromise commercial platforms for personal gain.